OpenStudy (anonymous):

Using the L'hopitale rule. What is Lim x^sinx x->0+

myininaya (myininaya):

\[e^{\ln(x^{\sin(x)})} =x^{\sin(x)} \text{ for } x>0 \\ e^{\ln(x^{\sin(x)})}=e^{\sin(x) \ln(x)}=e^\frac{\ln(x)}{\frac{1}{\sin(x)}}=e^{\frac{\ln(x)}{(\sin(x))^{-1}}}\]